# Mind the Value-Action Gap: Do LLMs Act in Alignment with Their Values?

### v1

@@ Appendix A Cultural and Social Values
We introduce the 56 universal values and their definitions outlined in the Schwartz’s Theory of Basic Values (Schwartz, 1994, 2012), which consists of 56 exemplary values covering ten motivational types. We show the complete list of value in Table 6.
@@ A Varied Life filled with challenge, novelty, and change Curious interested in everything, exploring
Table 6. The 56 universal values and their definitions outlined in the Schwartz’s Theory of Basic Values
@@ Appendix B Prompt Variation Design
Table 7. Human annotation performance on the eight prompts on data generation.
@@
Table 8. Human evaluation on the optimal two prompts with action feature attributions and natural language explanations.
@@
Table 9. Human evaluation for the generated data samples by annotators on Prolific from various countries.
@@ Appendix D Experiments of Predicting Actions with Explanations
Figure 6. The qualified prompt and examples.
@@ Appendix E More Findings
Figure 7. GPT4o-mini Model’s Heatmaps of (A) Task1, (B) Task2, and (C) Value-Action distance across 11 social topics.
@@
Figure 8. ChatGPT Model’s Heatmaps of (A) Task1, (B) Task2, and (C) Value-Action distance across 12 countries (left) and 11 social topics (right).
@@
Figure 9. Gemma2 Model’s Heatmaps of (A) Task1, (B) Task2, and (C) Value-Action distance across 12 countries (left) and 11 social topics (right).
@@
Figure 10. Llama3 Model’s Heatmaps of (A) Task1, (B) Task2, and (C) Value-Action distance across 12 countries (left) and 11 social topics (right).
@@
Figure 11. The GPT4o-mini’s results of ranking 56 values’ alignment distance on six countries: Australia, Canada, France, Egypt, India, Nigeria.
@@
Figure 12. The GPT4o-mini’s results of ranking 56 values’ alignment distance on six countries: Germany, United States, United Kingdom, Pakistan, Philippines, Uganda.
@@
Figure 13. The ChatGPT’s results of ranking 56 values’ alignment distance on six countries: Australia, Canada, France, Egypt, India, Nigeria.
@@
Figure 14. The ChatGPT’s results of ranking 56 values’ alignment distance on six countries: Germany, United States, United Kingdom, Pakistan, Philippines, Uganda.
@@
Figure 15. The Gemma2’s results of ranking 56 values’ alignment distance on six countries: Australia, Canada, France, Egypt, India, Nigeria.
@@
Figure 16. The Gemma2’s results of ranking 56 values’ alignment distance on six countries: Germany, United States, United Kingdom, Pakistan, Philippines, Uganda.
@@
Figure 17. The Llama3.3’s results of ranking 56 values’ alignment distance on six countries: Australia, Canada, France, Egypt, India, Nigeria.
@@
Figure 18. The Llama3.3’s results of ranking 56 values’ alignment distance on six countries: Germany, United States, United Kingdom, Pakistan, Philippines, Uganda.

### v2

@@ Appendix A Cultural and Social Values
We introduce the 56 universal values and their definitions outlined in the Schwartz’s Theory of Basic Values Schwartz (1994, 2012), which consists of 56 exemplary values covering ten motivational types. We show the complete list of value in Table 6.
@@ A Varied Life filled with challenge, novelty, and change Curious interested in everything, exploring
Table 6: The 56 universal values and their definitions outlined in the Schwartz’s Theory of Basic Values Schwartz (1992).
@@ Appendix B Prompt Variation Design
Table 7: Human annotation performance on the eight prompts on data generation.
@@
Table 8: Human evaluation on the optimal two prompts with action feature attributions and natural language explanations.
Metrics Definitions References
Correctness Whether the action accurately reflects agreement or disagreement with the stated value;  Bai et al. (2022)
Harmlessness Absence of harmful, offensive, or discriminatory content;  Bai et al. (2022)
Sufficiency Whether the action is sufficiently detailed to represent the value in the scenario;  DeYoung et al. (2019)
Plausibility Whether the action is realistic and feasible in the given situation.  Agarwal et al. (2024).
Table 9: The definition of evaluation metrics of human annotation process.
@@
Table 10: Human evaluation for the generated data samples by annotators on Prolific from various countries.
@@ Appendix D Experiments of Predicting Actions with Explanations
Figure 6: The qualified prompt and examples.
@@ Appendix E More Findings
Politics SocialNet Inequality Family Work Religion Env Identity Citizenship Leisure Health Sum
Deepseek 0.500 0.543 0.493 0.519 0.610 0.381 0.499 0.369 0.547 0.504 0.609 0.506
Qwen 0.365 0.468 0.299 0.395 0.406 0.373 0.316 0.273 0.373 0.386 0.484 0.376
Table 11: Averaged Value-Action Alignment Rates (i.e., F1 Scores) across 12 countries (top) and 11 social topics (bottom). The cell colors transition from bottom-2 through moderate to top-2 performances.
@@
Figure 7: GPT4o-mini Model’s Heatmaps of (A) Task1, (B) Task2, and (C) Value-Action distance across 11 social topics.
@@
Figure 8: ChatGPT Model’s Heatmaps of (A) Task1, (B) Task2, and (C) Value-Action distance across 12 countries (left) and 11 social topics (right).
@@
Figure 9: Gemma2 Model’s Heatmaps of (A) Task1, (B) Task2, and (C) Value-Action distance across 12 countries (left) and 11 social topics (right).
@@
Figure 10: Llama3 Model’s Heatmaps of (A) Task1, (B) Task2, and (C) Value-Action distance across 12 countries (left) and 11 social topics (right).
@@
Figure 11: The GPT4o-mini’s results of ranking 56 values’ alignment distance on six countries: Australia, Canada, France, Egypt, India, Nigeria.
@@
Figure 12: The GPT4o-mini’s results of ranking 56 values’ alignment distance on six countries: Germany, United States, United Kingdom, Pakistan, Philippines, Uganda.
@@
Figure 13: The ChatGPT’s results of ranking 56 values’ alignment distance on six countries: Australia, Canada, France, Egypt, India, Nigeria.
@@
Figure 14: The ChatGPT’s results of ranking 56 values’ alignment distance on six countries: Germany, United States, United Kingdom, Pakistan, Philippines, Uganda.
@@
Figure 15: The Gemma2’s results of ranking 56 values’ alignment distance on six countries: Australia, Canada, France, Egypt, India, Nigeria.
@@
Figure 16: The Gemma2’s results of ranking 56 values’ alignment distance on six countries: Germany, United States, United Kingdom, Pakistan, Philippines, Uganda.
@@
Figure 17: The Llama3.3’s results of ranking 56 values’ alignment distance on six countries: Australia, Canada, France, Egypt, India, Nigeria.
@@
Figure 18: The Llama3.3’s results of ranking 56 values’ alignment distance on six countries: Germany, United States, United Kingdom, Pakistan, Philippines, Uganda.
Appendix F Reasoned Explanations for Predicting Actions
We ground our approach in the Theory of Reasoned Action from social psychology Ajzen (1980); Fishbein and Ajzen (1980), which posits that identifying discrepancies between attitudes and behaviors is requisite to predict value-action gaps. Furthermore, we investigate whether reasoned explanations can aid in assessing the dynamics of value-action gaps in LLMs. To this end, we examine the reasoned explanations and highlighted action attributions included in the VIA dataset, and design a task to predict the alignment between value inclination and value-informed action. Concretely, we design a few-shot learning task where one observer model observes another target LLM’s contextual actions and explanations, and attempts to predict how the target LLM will state its value inclination given actions.
F.1 Explanations of Reasoning Actions Help Predict Value-Informed Actions
Act (baseline) Act+Attr Act+Exp Attr+Attr+Exp
GPT3.5-t 0.795 0.823 0.830 0.830
Llama3 0.778 0.797 0.823 0.820
Table 12: F1 scores of predicting the GPT4o-mini’s values based on only action or action with explanations and attributions.
In investigating how and to what extent value-action gaps can be predicted, we find that the inclusion of reasoned explanations improves the ability of an external model to predict the values of an LLM given their action selection. This yields a potential strategy for identifying and mitigating value-action gaps in real-world applications. For instance, when humans interact with LLMs in practical tasks, they can leverage reasoned explanations to guide LLMs toward value inclinations that align more closely with human expectations.
F.2 Risks in Value-Action Gaps
Category Level Risk Type Definition
Individual Discrimination Unequal treatment or representation based on race, gender, religion, disability, etc.
Autonomy Violation Manipulative or coercive suggestions that override individual agency.
Privacy Invasion Actions that cause distress, shame, anxiety, or erode self-worth.
Psychological Harm Disclosures or inferences that compromise personal data or identity.
Interaction Misleading Explanations Making inconsistent or misleading claims about its reasoning.
Overconfidence Presenting uncertain or incorrect actions with undue certainty.
User Manipulation Subtle steering of users toward actions that contradict their own values.
Societal Misinformation Spreading falsehoods, conspiracy, or misleading simplifications.
Polarization Amplifying societal divisions by aligning action with extreme or inconsistent stances.
Undermining Institutions Acting against values like justice or legality while claiming loyalty or fairness.
Table 13: The Definition and Value-Action Risk Taxonomy.